OpenStudy (anonymous):

1. Provide two polynomials and predict the number of complex roots for each. 2. You must explain how you found the number of complex roots for each.

OpenStudy (anonymous):

Please Help!!

OpenStudy (anonymous):

x^3=1 complex roots are omega & omega^2

OpenStudy (anonymous):

now how to find................

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

for any quadratic equation ax^2+bx+c=0 roots are x=\[[-b \pm \sqrt{b ^{2}-4ac}]/2a\]

OpenStudy (anonymous):

'Fundamental Theorem of Algebra'

OpenStudy (anonymous):

this b^2-4ac is called discriminent denoted by D if it is -ve then the roots are complex for eg x^2+x+1=0 you will get D=1-4=-3 so x=\[[-1\pm \sqrt{-3}]/2\] now \[\sqrt{-x}=i \sqrt{x}\] where i=square root of -1 so x=-1/2+i[3^(1/2)]/2 and other one both are complex:) do you want more on this

OpenStudy (anonymous):

That's all very interesting, @Aperogalics , but it does not answer the question asked by @katherinekc .

OpenStudy (anonymous):

oh sry i read it more concentratingly now:)

OpenStudy (anonymous):

yeah i have to put these two questions into an essay form

OpenStudy (anonymous):

paragraph*

OpenStudy (anonymous):

i have answer for that also:)

OpenStudy (anonymous):

@katherinekc type these phrases into a Google search: "Fundamental Theorem of Algebra" "Descartes' Rule of Signs"

OpenStudy (anonymous):

but what would two polynomials be? how would i predict the number of complex roots for each.

OpenStudy (anonymous):

You can make up the polynomials yourself. Whatever you want.

OpenStudy (anonymous):

and then the two phrases u gave would be how i found them?

OpenStudy (anonymous):

Research those phrases to get more background information. For example, I typed "Fundamental theorem of algebra" into Google and got: http://www.mathsisfun.com/algebra/fundamental-theorem-algebra.html Check it out.

OpenStudy (anonymous):

ok, thankyou

OpenStudy (anonymous):

simple way is to see the sign change in any polynomial as the no. of changes there are no. of zeros and if the degree of polynomial is greater by let n then there will be n complex roots:)

OpenStudy (anonymous):

@Aperogalics , you are not making any sense. Can you translate that into English, please?

OpenStudy (anonymous):

let me explain by an example :)........k